# Optimistic Thompson Sampling for No-Regret Learning in Unknown Games

### 07 Feb 2024 v1

@@ 5 Numerical Studies and Applications
Evaluation of the proposed algorithms on random matrix games and two real-world applications are conducted. IWE-Hedge (classical Exp3 algorithm) and IWE-RM serve as baseline comparisons, both adapted for bandit feedback scenarios. Moreover, Hedge and RM, which require full information, are utilized as well. Under the OTN learning framework (Algorithm 1), a total of four algorithms are compared with baselines: OTS-Hedge, OTS-RM, UCB-Hedge, and UCB-RM. Note that GP-MW (Sessa et al., 2019) is UCB-Hedge named here. Average expected regret is utilized as the performance metric. Appendix G.1 provides a detailed definition of the performance metrics and algorithm settings.
@@
(b) Best-resp. oppo.
(c) Stat.oppo.
Figure 3: Averaged regrets for different opponents on $50\times 50$ matrix game.
5.1 Random Matrix Games
This section evaluates different algorithms on two-player zero-sum matrix games. Each payoff matrix entry is an i.i.d. random variable generated from the uniform distribution $[-1,1]$, and each player has $M$ actions (payoff matrix is a squared matrix of size $M$). The total number of rounds is $T=10^{7}$. At each round, the two players receive noisy rewards ±$\pm\tilde{r}_{t}$, where $\tilde{r}_{t}=A_{ij}+\epsilon_{t}$ and $\epsilon_{t}\sim\mathcal{N}(0,0.1)$. Different matrix sizes are considered, i.e., $M=10,50,70,100$, and for each choice of $M$, $100$ independent simulation runs are conducted. The performance, averaged over $100$ simulation runs with different opponent models (see details in Appendix G.2) are compared.
Self-play
We first compare different algorithms under the self-play setting, where both players play the same algorithm. It can be observed from Fig.2(a) that algorithms exploiting the game structure outperform the two IWE baselines, especially when the matrix size is small. Among the algorithms, OTS-based algorithms converge faster than UCB-based algorithms. Furthermore, the average regret of RM decreases earlier than that of Hedge.
Against a best-response opponent
In this setting, a best-response opponent is employed, meaning the opponent is aware of the precise value of matrix $A$ and the player’s strategy at each round. Therefore, the opponent always chooses the best action that minimizes the expected payoff. The player continues to use the various algorithms described above. It can be observed from Fig.2(b) that all the algorithms in our proposed algorithmic framework outperform the IWE baselines. Once again, the OTS-based algorithms show a faster convergence behavior compared to the UCB-based algorithms.
Against a stationary opponent
A stationary opponent is considered, whose strategy is fixed to be a probability simplex over action space. It’s important to note that the average regret in such a situation indicates the efficiency on exploiting the weakness of the opponent. Convergence curves of average regret for $M=50$ are demonstrated in Fig. 2(c). The result clearly demonstrates that the OTS estimator brings considerable benefits in exploiting this weak opponent, compared to the IWE-based estimators.
Against an non-stationary opponent
The opponent’s strategy in this subsection becomes non-stationary, and the player can only hope a robust strategy for all possible opponent’s strategies. Specifically, a game matrix $A\in\mathcal{R}^{10\times 5}$ is generated, with each element sampling from $\mathcal{N}(0.5,2.0)$. The opponent’s actions are drawn from a fixed strategy that randomly changes every $50$ rounds. Each algorithm is evaluated for up to $1000$ rounds, and $100$ simulation runs are considered. Histograms of rewards over all rounds and simulation runs are plotted in Fig. 4. The percentage of negative rewards and mean reward values are summarized in Tab. 3. Compared with IWE, OTS has smaller percentage of negative rewards and achieves higher mean rewards. This indicates OTS is more robust than IWE.
5.2 Anti-jamming Problem: Linear Game
The anti-jamming problem is an important issue in signal processing literature (Song et al., 2011), which can be formulated as a non-cooperative game between a radar and a jammer.
@@ 5.2 Anti-jamming Problem: Linear Game
The competition takes place at the signal level, where both parties can modify the parameters of their transmitted signals. The radar’s objective is to avoid having its transmitted signal share the same carrier frequency as the jamming signal while the jammer seeks the opposite. In this experiment, we model this competition in the frequency domain as a linear game, with the signal-to-inference-plus-noise ratio (SINR) serving as the reward function $\text{SINR}(a,b;\theta)$. Further information on the anti-jamming game setting can be found in Appendix G.7.
We compare the average regret of different algorithms against an adaptive jammer, where the jammer takes action based on the radar’s latest $10$ actions (see details in Appendix G.7). As illustrated in Fig.5, the OTS-based and UCB-based algorithms outperform the IWE-based algorithms.
@@
Figure 5: Averaged regrets in the anti-jamming problem.
@@ 5.3 Repeated Traffic Routing: Kernelized Game
In this subsection, we consider the traffic routing problem from the transportation literature, which can be modeled as a multi-player game over a directed graph. Each node pair in the graph represents an individual player, and each player $i$ seeks to find the best route to send $U^{i}$ units from its origin node to its destination node. The travel time serves as the reward, depending on the traversed edges’ total occupancy. Each player’s action set is the available route in the graph, and the negative travel time of the route is the reward. A detailed description of the problem setup is provided in Appendix G.8.
We use the Sioux-Falls road network dataset (Bar-Gera, 2015) in our experiment. This network is a directed graph with $24$ nodes and $76$ edges, and there are a total of $N=528$ players. Each player $i$’s action space $\mathcal{A}^{i}$ is specified by the $5$ shortest routes, and any route that is more than three times longer than the shortest route is removed from $\mathcal{A}^{i}$. To capture the correlations in the reward function, Gaussian Process is utilized to estimate the reward function (Sessa et al., 2019). Details of the reward function estimation can be found in Appendix G.8. It is worth noting that GP-MW proposed in (Sessa et al., 2019) is equivalent to UCB-Hedge in this paper. In addition to regret, we also consider average congestion (Sessa et al., 2019) as a performance metric. Fig. 6 demonstrates that OTS-based and UCB-based algorithms outperform the IWE-based algorithms. The three proposed new algorithms, OTS-Hedge, OTS-RM, and UCB-RM, outperform UCB-Hedge (Sessa et al., 2019).
Figure 6: Average regret and congestion in traffic routing problem.
6 Conclusion
In this study, we introduced Thompson Sampling (TS)-type algorithms that utilize insights into opponents’ actions and reward structures, significantly optimizing experimental budgets with over a tenfold reduction relative to standard algorithms. Additionally, we developed the Optimism-then-NoRegret (OTN) learning framework, a versatile tool for navigating the complexities of unknown games, which encompasses a range of game algorithms as specific instances. The efficacy of our proposed solutions is validated through substantial performance enhancements observed in both synthetic scenarios and real-world applications.
Looking ahead, the convergence of emerging techniques that approximate Thompson Sampling within complex models (Li et al., 2022, 2023), such as deep neural networks, presents a fertile ground for innovation. The integration of these approaches with the OTN framework and the Optimistic Thompson Sampling (OTS) algorithm, particularly in multi-agent environments of increased complexity, holds considerable promise. Such advancements could further refine decision-making processes, enhance model robustness, and expand the applicability of our methodologies to a broader spectrum of challenges in machine learning and artificial intelligence. This prospective synergy between advanced sampling techniques and the OTN framework is poised to push the boundaries of what’s achievable in the realm of machine learning, offering exciting possibilities for future research and application.
@@ Appendix A Additional discussion on related works
Appendix B Description of Games and Full Information Game
@@
Table 4: Game Examples.
@@ B.2 Full Information Feedback
Appendix C Updating Rules of Different Games
@@ Appendix D Details for Importance Weighted Estimator in Section 3
6:     Update $\hat{X}_{t+1}$ and $X_{t+1}$ with Equations 11 and 12 and $\gamma_{t}$.
@@
Suppose at round $t$, Alice choose $A_{t}\sim P_{X_{t}}$ and receive the feedback $R_{t+1,A_{t},B_{t}}$. By algorithm 3 and Equations 10, 11 and 12, If $\sum_{a}\tilde{C}_{t-1,a}^{+}\leqslant 0$, then obviously $\tilde{C}^{+}_{t-1,a}=0$ for all action $a\in\mathcal{A}$. Then, the lemma trivially holds.
@@
Appendix E Analysis of the failure for Thompson sampling estimator in Section 3.3
@@ E.2 TS-RM suffers linear regret
E.3 Why Optimistic variant of TS can converge?
@@ G.8 Repeated traffic routing problem
where $k_{\rm L}(\cdot,\cdot)$ and $k_{\rm P}(\cdot,\cdot)$ are linear and polynomial kernels respectively. The hyperparameters of kernels are set the same as in (Sessa et al., 2019). Details on GP for estimating reward functions can be found in Section 2 of (Sessa et al., 2019) and we do not repeated them here666 The implementation refers the code released by authors of (Sessa et al., 2019) at https://github.com/sessap/stackelucb. Except the regret, the congestion of edge $e$ is also included as a performance metric for a joint action $\bm{a}$, which is computed as $0.15(\sum_{i=1}^{N}[a^{i}]_{e}/C_{e})^{4}$. The averaged congestion is obtained by averaging congestion over all edges.

### 25 Feb 2024 v2

@@ 5 Numerical studies and applications
We conduct evaluations of the proposed algorithms on random matrix games and two real-world applications. IWE-Hedge (the classical Exp3 algorithm) and IWE-RM are used as baseline comparisons, both adapted for bandit feedback scenarios. Additionally, Hedge and RM, which necessitate full information, are also employed. Within the OTN learning framework (Algorithm 1), four algorithms: OTS-Hedge, OTS-RM, UCB-Hedge, and UCB-RM, are compared against these baselines. It is noteworthy that GP-MW (Sessa et al., 2019) is referred to as UCB-Hedge in this context. The performance metric used is the average expected regret, with a detailed definition of performance metrics and algorithm settings provided in Appendix G.1.
5.1 Random Matrix Games
We assess various algorithms in repeated two-player zero-sum matrix games, where each entry of the payoff matrix is an independently and identically distributed random variable from the uniform distribution $[-1,1]$. Players have $M$ actions each, forming a square payoff matrix of size $M$. The experiment is run for a total of $T=10^{7}$ rounds, during which players receive noisy rewards ±$\pm\tilde{r}_{t}$, with $\tilde{r}_{t}=A_{ij}+\epsilon_{t}$ and $\epsilon_{t}\sim\mathcal{N}(0,0.1)$. The analysis covers different matrix sizes ($M=10,50,70,100$), with each scenario tested across $100$ independent simulation runs. Performance is averaged over these runs against varying opponent models, detailed further in Appendix G.2.
@@ 5.1 Random Matrix Games
(b) Best-response opponent
(c) Stationary opponent
Figure 3: Averaged regrets for different opponents in a $50\times 50$ matrix game.
Self-play.
In the self-play scenario, where both players employ the same algorithm, Figure 2(a) reveals that algorithms leveraging the game’s structure significantly outperform the IWE baselines, particularly in smaller matrix sizes. Notably, OTS-based algorithms demonstrate quicker convergence than their UCB-based counterparts, with RM algorithms showing an earlier reduction in average regret compared to Hedge algorithms.
Best-response opponent.
This scenario introduces a best-response opponent, fully informed about the matrix $A$ and the player’s current strategy, hence always opting for the action that minimizes the player’s expected payoff. Figure 2(b) shows that all algorithms within our proposed framework surpass the performance of the IWE baselines, with OTS-based algorithms again displaying superior convergence speeds relative to UCB-based methods.
Stationary opponent.
The focus here is on a stationary opponent whose strategy is a fixed probability distribution over the action space. The average regret in this context is indicative of an algorithm’s effectiveness at exploiting the opponent’s static behavior. As depicted in Figure 2(c), OTS estimators confer significant advantages in exploiting such opponents over IWE-based estimators, showcasing the efficacy of leveraging game structure in algorithm design.
Non-stationary opponent.
Contrasting with previous settings, the opponent’s strategy here varies non-statically, altering every $50$ rounds based on a predefined pattern. The game matrix $A\in\mathcal{R}^{10\times 5}$, with entries drawn from $\mathcal{N}(0.5,2.0)$, presents a challenging dynamic environment. The algorithms’ robustness is evaluated over $1000$ rounds across $100$ simulation runs, with the rewards’ distribution and performance metrics summarized in Figure 4 and Table 2, respectively. The OTS algorithms notably yield fewer negative rewards and higher average rewards than their IWE counterparts, illustrating enhanced resilience in face of strategic variability.
@@
Figure 4: Reward histograms of different algorithms against a non-stationary opponent, illustrating the performance disparity between algorithm types. OTS-based and UCB-based algorithms demonstrate significant efficiency, particularly in adapting to the jammer’s changing strategies.
5.2 Two-player radar signal processing: Linear game
The anti-jamming problem, a critical issue in signal processing literature (Song et al., 2011), is modeled as a non-cooperative game between a radar and a jammer. The strategic interaction at the signal level involves both parties adjusting their transmitted signals’ parameters to achieve opposing objectives: the radar aims to avoid signal interference by differing its carrier frequency from the jammer’s, while the jammer attempts to match it. This competition is formalized in the frequency domain as a linear game, with the signal-to-interference-plus-noise ratio (SINR) serving as the reward function $\text{SINR}(a,b;\theta)$. We conduct experiments comparing the average regret of various algorithms against an adaptive jammer, which updates its action based on the radar’s most recent $10$ actions. As depicted in Fig.5, OTS-based and UCB-based algorithms significantly outperform IWE-based algorithms, underscoring the advantages of adaptive strategies in anti-jamming scenarios.
@@ 5.2 Two-player radar signal processing: Linear game
Figure 5: Averaged regrets in the anti-jamming problem demonstrate the superior efficiency of OTS-based and UCB-based algorithms over IWE-based ones. Notably, OTS-Hedge requires significantly fewer samples to achieve a comparable level of average regret, illustrating its robustness and adaptability. Specifically, to achieve an average regret of $10^{-2}$, OTS-Hedge requires only $3.7\%$ and $31\%$ of the samples needed by IWE-Hedge and UCB-Hedge, respectively; OTS-RM uses only 0.1% of IWE-RM’s, 25% of UCB-RM’s, and 53% of OTS-Hedge’s samples.
Further details on the anti-jamming game setting, including experimental setup and algorithm parameters, are provided in Section G.7, offering readers a comprehensive understanding of the methodologies employed to achieve these results.
@@ 5.3 Repeated Traffic Routing: Kernelized Game
The traffic routing problem, as derived from transportation studies, is formulated as a multi-agent game on a directed graph. In this model, each node pair signifies a distinct player, tasked with routing $U^{i}$ units from an origin to a destination node. The objective is to minimize travel time, influenced by the cumulative occupancy of traversed edges, making it the reward function. The set of actions available to each player comprises all feasible routes within the graph, with rewards inversely proportional to travel times. For an in-depth exploration of the setup, refer to Section G.8.
Utilizing the Sioux-Falls road network dataset (Bar-Gera, 2015), we simulate a network of $24$ nodes and $76$ edges, engaging $N=528$ players. The action space $\mathcal{A}^{i}$ for player $i$ includes up to the $5$ shortest paths, excluding any route exceeding triple the shortest path’s length. A Gaussian process models the reward function’s correlations, as detailed in Section G.8, following the methodology of GP-MW (Sessa et al., 2019), here referred to as UCB-Hedge. Performance metrics extend beyond regret to encompass average congestion, offering a holistic view of traffic dynamics (Sessa et al., 2019). As depicted in Figure 6, OTS and UCB algorithms surpass IWE counterparts, highlighting the efficacy of the newly introduced OTS-Hedge, OTS-RM, and UCB-RM algorithms in surpassing UCB-Hedge’s performance.
Figure 6: Illustrating average regret and congestion for the traffic routing challenge. Notably, OTS-RM showcases unparalleled efficiency, requiring significantly fewer samples than its counterparts to achieve benchmark regret and congestion levels, thereby underscoring its capacity to manage hundreds of agents with reduced sample complexity. Specifically, to achieve an average regret of $750$, OTS-RM requires $8\%$ of IWE-Hedge’s, $16.3\%$ of UCB-Hedge’s, $21.6\%$ of OTS-Hedge’s, and $72.7\%$ of UCB-RM’s samples. For an average congestion of $10$, OTS-RM uses $5.6\%$ of IWE-Hedge’s, $15.3\%$ of UCB-Hedge’s, $20\%$ of OTS-Hedge’s, and $66.7\%$ of UCB-RM’s samples. Moreover, OTS-RM and can effective handle hundreds of decision-makers in the most sample-efficient manner.
6 Conclusions
This study presents a breakthrough in leveraging opponents’ actions and reward structures through Thompson Sampling (TS)-inspired algorithms, which markedly optimize experimental resources, reducing costs by more than tenfold relative to traditional methods. We also introduce the Optimism-then-NoRegret (OTN) learning framework, an adaptive strategy for mastering the intricacies of unknown games, which subsumes various algorithmic approaches as special cases. Our proposed techniques have demonstrated significant performance improvements in both simulated and real-world settings. Future directions point towards integrating emerging TS approximation methods within complex models(Li et al., 2022, 2023, 2024), like deep neural networks, with the OTN framework and the Optimistic Thompson Sampling (OTS) algorithm, especially in more complex multi-agent environments. This convergence is expected to refine decision-making processes, bolster model robustness, and widen the scope of our methodologies to encompass a greater array of machine learning and artificial intelligence challenges. The anticipated fusion of advanced sampling techniques with the OTN framework heralds a new frontier in machine learning research, brimming with promising prospects for innovation and application.
@@ 6 Conclusions
Appendix: Optimistic Thompson Sampling for No-Regret Learning in Unknown Games
@@ Appendix A Additional discussion on related works
Appendix B Description of bandit games and full information game
@@
Table 3: Examples of bandit games.
@@ B.2 Full information feedback
Appendix C Updating rules in various games
@@ Appendix D Details for importance weighted estimator in Section 3
6:     Update $\hat{X}_{t+1}$ and $X_{t+1}$ with Equations 8 and 9 and $\gamma_{t}$.
@@
Suppose at round $t$, Alice choose $A_{t}\sim P_{X_{t}}$ and receive the feedback $R_{t+1,A_{t},B_{t}}$. By algorithm 3 and Equations 7, 8 and 9, If $\sum_{a}\tilde{C}_{t-1,a}^{+}\leqslant 0$, then obviously $\tilde{C}^{+}_{t-1,a}=0$ for all action $a\in\mathcal{A}$. Then, the lemma trivially holds.
@@
Appendix E Failure analysis for Thompson sampling in Section 3.3
@@ E.2 TS-RM suffers linear regret
E.3 Why optimistic variant of TS would not suffer linear regret?
@@ G.8 Repeated traffic routing problem
where $k_{\rm L}(\cdot,\cdot)$ and $k_{\rm P}(\cdot,\cdot)$ are linear and polynomial kernels respectively. The hyperparameters of kernels are set the same as in (Sessa et al., 2019). Details on GP for estimating reward functions can be found in Section 2 of (Sessa et al., 2019) and we do not repeated them here333 The implementation refers the code released by authors of (Sessa et al., 2019) at https://github.com/sessap/stackelucb. Except the regret, the congestion of edge $e$ is also included as a performance metric for a joint action $\bm{a}$, which is computed as $0.15(\sum_{i=1}^{N}[a^{i}]_{e}/C_{e})^{4}$. The averaged congestion is obtained by averaging congestion over all edges.